OpenStudy (anonymous):

If f(x)=ln(9+8x) find f'(x) . How do I find the derivative with ln?

OpenStudy (bahrom7893):

\[(ln(f(x)))' = \frac{1}{f(x)}*f'(x)\]

OpenStudy (anonymous):

so would it be \[\frac{ 1 }{ \ln(9+8x)} *\ln(9+8x)\] at first?

OpenStudy (bahrom7893):

No, you have: \[f(x)=9+8x\]

OpenStudy (anonymous):

as an answer?

OpenStudy (bahrom7893):

You need to find the derivative of: \[ln(f(x))\]